OpenStudy (anonymous):

Need help a chess competition has eliminations each round. The table below shows the number of players in each of the first 5 rounds of tournament Round(x) Players f(x) 1 256 2 128 3 64 4 32 5 16 Compute the average rate of change of f(x) from x = 1 to x = 5 identify the meaning of that rate.

Vocaloid (vocaloid):

[ f(5)-f(1) ] / (5-1) find f(5) and f(1), then plug them into the equation

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

I'm terrible at math can explain how to do the problem

Vocaloid (vocaloid):

what is f(5)? (look at the table)

OpenStudy (anonymous):

16?

Vocaloid (vocaloid):

yeah! now find f(1)

OpenStudy (anonymous):

256?

Vocaloid (vocaloid):

yup! now plug those values into our expression [f(5)-f(1)] / (5-1)

OpenStudy (anonymous):

F(16)-f(256)/ (5-1)?

Vocaloid (vocaloid):

uh, not quite, the f's go away when you replace them (16-256) / (5-1) = ?

OpenStudy (anonymous):

-60 would be the answer?

Vocaloid (vocaloid):

yup!

OpenStudy (anonymous):

So it would be -60; on average, there was a loss of 60 each round?

Vocaloid (vocaloid):

yeah! you got it

OpenStudy (anonymous):

Thank you! Can you help me with a few more?

Vocaloid (vocaloid):

sure

OpenStudy (anonymous):

Ok which of the following options results in a graph that shows exponential growth? f(x) = 0.4(0.2)^x f(x) = 4(0.98)^x f(x) = 0.7(5)^x f(x) = 5(4)^-x

Vocaloid (vocaloid):

2 requirements for exponential growth: 1. the exponent must be positive 2. the ratio (the number in parentheses) must be greater than 1 given this information, which of the four choices do you think it is?

OpenStudy (anonymous):

F(x) = 0.7(5)^x?

Vocaloid (vocaloid):

yeah, you got it ~

OpenStudy (anonymous):

Ok 3 more

OpenStudy (anonymous):

Which of the following options is an equivalent function to f(x) = 4(3)^2x f(x) = 36^x f(x) = 4(9)^x f(x) = 144^x f(x) = 4(6x)

Vocaloid (vocaloid):

you can simplify (3^2)^x to 9^x, but you have to leave the 4 alone since it's not attached to the x f(x) = 4(9)^x

OpenStudy (anonymous):

Wait f(x) = 4(9)^x is the answer?

Vocaloid (vocaloid):

yes

OpenStudy (anonymous):

Oh ok I would say explain but it seems you already did ok 2 more

OpenStudy (anonymous):

For f(x) = 0.01(2)^x, find the average rate of change from x = 12 and x = 15 3 40.96 95.57 286.72

Vocaloid (vocaloid):

it's pretty much the same as the first problem you showed me [ f(15)-f(12) ] / ( 15-12) = ?

OpenStudy (anonymous):

I got 1 but it's not right

Vocaloid (vocaloid):

ok, let's take it step by step f(15) = ?

OpenStudy (anonymous):

x?

Vocaloid (vocaloid):

f(x) = 0.01(2)^x f(15) = ?

OpenStudy (anonymous):

I'm confused I'm so bad at math so your plugging in the 15 so it's f(15) = 0.01(2)^x?

Vocaloid (vocaloid):

when you plug in 15, you replace "x" with 15 so f(15) = 0.01(2)^15

OpenStudy (anonymous):

Oh ok

Vocaloid (vocaloid):

so, f(15) = ?

OpenStudy (anonymous):

So what do you do with f(15) = 0.01(2)^15

OpenStudy (anonymous):

0.01(2)^15

Vocaloid (vocaloid):

just type it out on a calculator and tell me what you get

OpenStudy (anonymous):

I don't have that calculator I have to find it

Vocaloid (vocaloid):

you can use an online calculator

OpenStudy (anonymous):

327.68

Vocaloid (vocaloid):

right, now find f(12)

OpenStudy (anonymous):

40.96

Vocaloid (vocaloid):

right, now go back to the formula I gave you back then [f(15)-f(12)] / (15-12) = ?

OpenStudy (anonymous):

286.72

OpenStudy (anonymous):

?

Vocaloid (vocaloid):

you still need to divide by (15-12)

OpenStudy (anonymous):

I did I divided 286.72 by 286.72 and I got 1 or do you divide 286.72 by 3

Vocaloid (vocaloid):

you divide by 3

OpenStudy (anonymous):

95.57?

Vocaloid (vocaloid):

yeah! that's our answer

OpenStudy (anonymous):

Ok this ones a graph

OpenStudy (anonymous):

Identify the domain of the exponential function shown in the following graph: All real numbers All positive numbers 1 less than or equal to x less than or equal to 30 All integers

OpenStudy (anonymous):

Idk how to to add the graph I'm on my phone

Vocaloid (vocaloid):

can you try to draw what it looks like?

OpenStudy (anonymous):

How do you draw it?

Vocaloid (vocaloid):

I'm not sure what you would do on a phone

Vocaloid (vocaloid):

Is there a "draw" button anywhere?

OpenStudy (anonymous):

Ok hold on

OpenStudy (anonymous):

A curve runs just above the x axis through the second quadrant, passes through point 0 comma 1 and continues off screen near 1 comma 12.

Vocaloid (vocaloid):

it seems like the domain would be all real values since there aren't any gaps or asymptotes

OpenStudy (anonymous):

Oh ok

Vocaloid (vocaloid):

the first one is the best choice here

OpenStudy (anonymous):

I can try attach the graph if it would help strengthen the answer

OpenStudy (anonymous):

No need I got them all right thank you!!!

Vocaloid (vocaloid):

great